# Reinforced MLLM: A Survey on RL-Based Reasoning in Multimodal Large Language Models

### v1

@@ 4. RL-Based MLLM Reasoning
R1-Omni (Zhao et al., 2025) HumanOmni-0.5B SFT&RL GRPO Emotion T
Table 2. Overview of multimodal reasoning models employing reinforcement learning strategies. Univ. Reason denotes general-purpose reasoning across diverse tasks, Image Quality refers to the image quality understanding task, Emotion represents the emotion recognition task. In the MM-CoT column, T indicates support for the text modality, I for the image modality.
Model Reward
KIMI K1.5 (Team et al., 2025) $R=r_{acc}(x,y,y^{*})\in\{0,1\}$
MedVLM-R1 (Pan et al., 2025) $R=r_{acc}+r_{format}$
Visual-RFT (Liu et al., 2025d)
Detection Tasks: $R=r_{IoU}+r_{conf}+r_{format}$
Classification Task:$R=r_{acc}+r_{format}$
R1-Omni (Zhao et al., 2025) $R=r_{acc}+r_{format}$
VisualThinker-R1-Zero (Zhou et al., 2025) $R=r_{acc}+r_{format}$
Vision-R1 (Huang et al., 2025) $R=\left\{\begin{matrix}1,&\text{Acc}\text{ and }\text{Format}\\ 0,&\text{otherwise}\end{matrix}\right.$
Seg-Zero (Liu et al., 2025c) $R=r_{BIoU}+r_{BL1}+r_{PL1}+r_{SFormat}+r_{format}$
MM-Eureka-8B (Meng et al., 2025) $R=r_{acc}+\lambda r_{format}$
MM-Eureka-Zero-38B (Meng et al., 2025) $R=r_{acc}+\lambda r_{format}$
LMM-R1 (Peng et al., 2025) $R=r_{acc}+\lambda r_{format}$
R1-Onevision (Yang et al., 2025) $R=r_{acc}+\lambda r_{format}$
R1-AQA(Li et al., 2025a) $R=r_{acc}+\lambda r_{format}$
R1-VL (Zhang et al., 2025) $r_{auc}^{i}(s_{t},a_{t},s_{t+1})=\begin{cases}1+\alpha k^{i},&\mathrm{ans}(s_{t+1})=y\\ \alpha k^{i},&\mathrm{ans}(s_{t+1})\neq\mathrm{null},\neq y\\ 0,&\mathrm{ans}(s_{t+1})=\mathrm{null}\end{cases},$ $r_{val}^{i}(s_{t},a_{t},s_{t+1})=\begin{cases}1,&\mathbb{I}(\delta^{c}(s_{t+1}))\cdot\mathbb{I}(\delta^{l}(s_{t+1}))=1\\ 0,&\text{otherwise}\end{cases},$ $R=r_{auc}+r_{val}$
Skywork R1V (Yi Peng, 2025) $R=r_{acc}+\lambda r_{format}$
OThink-MR1 (Liu et al., 2025e) $R=r_{acc}+\lambda r_{format}$
OpenVLThinker-7B (Deng et al., 2025a) $R=r_{acc}$
UI-R1 (Lu et al., 2025) $R=R_{\tau}+R_{C}+R_{\mathcal{F}},$ $\tau:\text{action type},$ $C:\text{clock coordinate },$ $\mathcal{F}:\text{format}$
Reason-RFT (Tan et al., 2025) $R=r_{acc}+r_{format}$
Q-Insight (Li et al., 2025b) $r^{(i)}=r^{(i)}_{\text{fmt}}+\mathbf{1}_{\text{scr}}\cdot r^{(i)}_{\text{scr}}+\mathbf{1}_{\text{deg}}\cdot\left(\alpha_{1}\cdot r^{(i)}_{\text{deg}}+\alpha_{2}\cdot r^{(i)}_{\text{lev}}\right)$
Table 3. Reward function for ecah MLLM RL-based reasoning method.
Deepseek-R1 (Guo et al., 2025) demonstrates that the reasoning capabilities of LLMs can be activated through pure RL. Currently, RL-based reasoning methods model the reasoning process as a Markov Decision Process (MDP). The core objective is to enhance the model’s reasoning capability by maximizing the expected reward of the reasoning path $S$. Inspired by the success of Deepseek-R1, researchers have begun exploring how to apply the R1 training paradigm to MLLMs to further improve their reasoning performance in complex tasks. In Table 2, we summarize the current research progress of reasoning methods based on reinforcement learning in the multimodal domain. Additionally, in Table 3, we provide a detailed list of the reward function designs corresponding to each method, helping readers better understand the design principles behind different approaches.
In Section 4.1, we systematically review recent attempts to adapt the R1 training paradigm from LLMs to MLLMs and analyze the key challenges and solutions involved. Subsequently, Section 4.2 focuses on the design of different reward mechanisms in multimodal scenarios, categorizing them into two types: Outcome-supervised Reward Mechanism and Process-supervised Reward Mechanism. Section 4.3 focuses on the efforts of different methods to leverage training data for improving training efficiency and stability. Finally, Section 4.4 summarizes the currently popular RLHF frameworks, which can facilitate the rapid iteration of RL-based reasoning methods.
@@ 4.1.1. Standardized R1 Training Paradigm in LLM
The success of Kimi K1.5 (Team et al., 2025) and DeepSeek R1 (Guo et al., 2025) in enhancing model reasoning capabilities through reinforcement learning has driven the rise of the R1 training paradigm in the field of model reasoning.
@@
Kimi K1.5 (Team et al., 2025) employs a variant of the online policy mirror descent (OPMD) algorithm (Abbasi-Yadkori et al., 2019; Mei et al., 2019; Tomar et al., 2020) to apply reinforcement learning to MLLMs, enhancing their reasoning capabilities in multimodal domains. Additionally, it adopts a length penalty mechanism to prevent the model from overthinking. This represents a successful attempt in the direction of MLLM reasoning based on RL. Meanwhile, DeepSeek R1 (Guo et al., 2025) introduces a key innovation by validating that Reinforcement Learning with Verifiable Reward (RLVR) is sufficient to effectively elicit reasoning capabilities in LLMs. The core of RLVR lies in its rule-driven reward mechanism, which optimizes model performance efficiently and reliably. Compared to traditional SFT and process supervision methods, RLVR demonstrates significant advantages when handling verifiable tasks, such as mathematical problem-solving and code generation: it eliminates the need for labor-intensive data generation and annotation while significantly improving model performance with limited training data. Additionally, by avoiding the complex training of reward models, RLVR effectively mitigates the issue of reward hacking.
@@
(12) $R(q,o)=\begin{cases}1,&\text{if }o=\text{ground truth},\\ 0,&\text{otherwise}.\end{cases}$
@@
where $R_{format}$ evaluates whether the output of the policy model $\pi_{\theta}$ adheres to predefined template specifications. For example, the template may require the output to follow the format ¡think¿¡/think¿¡answer¿¡/answer¿, where ¡think¿¡/think¿ contains the model’s reasoning process, and ¡answer¿¡/answer¿ includes the final result. On the other hand, $R_{acc}$ focuses on assessing whether the output in the ¡answer¿¡/answer¿ section matches the ground truth. Through this design, RLVR not only ensures the accuracy of the model’s output but also guides the model to generate structured and interpretable reasoning content, thereby enhancing both performance and explainability. This approach has proven to be highly efficient and reliable in practice, particularly for complex tasks.
@@
While RL has shown promise in improving the reasoning abilities of large language models, relying solely on large-scale RL is insufficient for robust reasoning development. One fundamental issue is that RL algorithms typically require a reasonably good initialization to be effective: without a meaningful starting policy, exploration becomes inefficient, rewards are too sparse, and training often collapses to trivial or degenerate behaviors. In reasoning tasks, where the action space is combinatorially large and intermediate rewards are hard to define, a model starting from random or weak behavior struggles to discover successful reasoning strategies purely through trial and error. Cold start methods aim to address this gap by providing initial scaffolding — through techniques such as curriculum design, synthetic demonstrations, weak supervision, or structured exploration — to bootstrap the model into a regime where RL can be applied effectively. DeepSeek R1 collected thousands of cold-start data points to fine-tune the base model, enhancing the readability of reasoning CoT and the potential for RL. Without such cold start mechanisms, scaling up RL alone often leads to prohibitively high sample complexity, instability, and limited generalization across reasoning tasks. Thus, successful reasoning training typically involves a hybrid approach: cold start strategies to seed useful reasoning behaviors, followed by reinforcement learning to refine and optimize them over time.
@@ 4.1.2. R1 Training Paradim for MLLM
Inspired by the design of format reward and accuracy reward mechanisms in Equation 13 within DeepSeek R1-Zero, researchers began attempting to transfer this successful experience to multimodal tasks. MedVLM-R1 (Pan et al., 2025) extended the DeepSeek R1 training paradigm to visual question-answering (VQA) tasks in the medical domain, guiding the model to answer medical multiple-choice questions through explicit reasoning paths, thereby improving prediction accuracy and out-of-distribution (OOD) generalization performance. Similarly, VisualThinker-R1-Zero (Zhou et al., 2025) and MM-Eureka (Meng et al., 2025) successfully replicated DeepSeek R1’s long CoT characteristics, including the ”Aha Moment,” self-reflection mechanisms, and dynamic adjustment of response length, in spatial reasoning and mathematical VQA tasks, respectively. These features demonstrate that the experience of DeepSeek R1 can be effectively transferred to complex multimodal reasoning tasks.
OThink-MR1 (Liu et al., 2025e) improved the widely used GRPO algorithm by introducing dynamic weighting for the KL divergence term, balancing exploration in the early training stages with exploitation in later stages. In the early stages, reducing the weight of the KL divergence term allows for more thorough exploration of the solution space, while in later stages, increasing the weight aligns the model with the reference model for better performance. Despite these efforts achieving the application of the R1 training paradigm in multimodal tasks, they remain constrained by Equation 13 format rewards and accuracy rewards, somewhat neglecting the unique characteristics of tasks in the multimodal domain.
@@
To improve the stability of the RL training process and the upper limit of the model’s inference performance, some methods have begun constructing multi-modal SFT data to cold-start the base model. Vision-R1 (Huang et al., 2025) first leveraged VLM and DeepSeek R1 to construct a batch of reasoning cold-start data, enabling VLM to initially acquire reasoning capabilities. Subsequently, through Progressive Thinking Suppression Training (PTST)—a stage-wise RL strategy that progressively extends CoT length while separating format and accuracy rewards—it decoupled format reward and accuracy reward at different stages of RL training, gradually increasing CoT length from short to long and alleviating overthinking from the outset. Due to the relative scarcity of image-text data compared to pure text data, LMM-R1 (Peng et al., 2025) proposed a two-stage training strategy to make efficient use of textual CoT data: the first stage involves RL training on pure text data to enhance MLLM’s foundational text reasoning abilities, while the second stage incorporates image-text data to further improve the model’s generalization capabilities in visual perception and other multimodal tasks. R1-Onevision (Yang et al., 2025) points out that an increase in the length of reasoning paths may cause the model to overlook visual information. To enable MLLMs to actively integrate multimodal information during reasoning, the method divides the training process into two stages: SFT and RL. In the SFT stage, R1-Onevision constructs a CoT-based reasoning dataset, which includes formalized image descriptions. By standardizing the model’s reasoning paths, it ensures that these paths incorporate formalized image descriptions. In the RL stage, the model’s reasoning capabilities and OOD generalization are further strengthened. Skywork R1V (Yi Peng, 2025) proposes a lightweight R1 paradigm training method to alleviate the high consumption of hardware resources. Unlike full fine-tuning methods, this approach combines a vision encoder with an LLM backbone possessing reasoning capabilities via a lightweight visual projector module. Iterative training of the visual projector module using SFT and the R1 paradigm progressively enhances MLLM’s reasoning performance in visual tasks. To address the diversity of mathematical problems, Reason-RFT has designed customized reward mechanisms for discrete-value problems, mathematical problems, and function-based problems. Each mechanism is tailored to the unique characteristics of its respective problem category, ensuring precise and fair evaluations.
@@
Beyond image-text interactions, some research has explored applying the R1 training paradigm to tasks involving other modalities. For instance, R1-Omni (Zhao et al., 2025) conducted preliminary explorations in mixed-modality tasks involving video and audio. It first performs SFT cold-start with a small amount of reasoning data, then utilized the R1 training paradigm to explore the application of reasoning models in emotion recognition tasks. Additionally, R1-AQA (Li et al., 2025a) undertook a series of reinforcement learning-related explorations in audio understanding and reasoning tasks, providing new insights for the further development of multimodal reasoning.
@@ 4.2.1. Outcome-supervised reward Mechanism
Outcome-supervised reward mechanisms represent an advancement in aligning RL with complex, multimodal tasks by leveraging task-specific and cross-modal feedback. These mechanisms not only enhance model performance through structured and hierarchical rewards but also pave the way for more interpretable and generalizable reasoning across diverse modalities.
@@
Given the complexity and diversity of multimodal tasks, recent studies have explored reward designs that leverage the intrinsic properties of each modality to broaden the applicability of RL. Visual-RFT (Liu et al., 2025d) targets fundamental computer vision problems such as image classification, object detection, and image-text grounding, introducing task-specific visual rewards: classification tasks use label matching as the reward signal; detection tasks optimize IoU between predicted and ground-truth boxes; and grounding tasks evaluate alignment between generated text and image regions. These rewards are all based on final prediction accuracy, providing explicit label-supervised signals in the visual modality. However, these methods still design rewards within a single modality, without incorporating cross-modal interactions, thus representing a transitional stage in the broader shift from unimodal to multimodal reward modeling.
@@
To address the non-linear interactions in multimodal reasoning, recent studies assign modality-specific rewards to promote more active cross-modal interactions. UI-R1 (Lu et al., 2025) introduces a multimodal reward framework into GUI-based interaction tasks by jointly evaluating the predicted action type (e.g., click or scroll), action arguments (e.g., position and target object), and output formatting validity (e.g., JSON schema). This reward mechanism is effectively combining visual input and textual instructions, establishing an initial form of alignment feedback between the two modalities. Q-Insight (Li et al., 2025b) follows a similar direction in the context of image quality assessment by jointly optimizing score regression and degradation perception through GRPO-based training. Its reward design integrates verifiable score accuracy with distortion classification, enabling the model to perform perceptual reasoning with minimal supervision and strong generalization to zero-shot comparison tasks. Seg-Zero (Liu et al., 2025c) extends this idea to the field of visual segmentation. In addition to pixel-wise IoU and L1 distance metrics for segmentation quality, Seg-Zero (Liu et al., 2025c) introduces model-generated CoT responses as part of the output, enabling partial supervision over the reasoning trace. Although the reward still targets final segmentation accuracy, the presence of CoT outputs begins to connect perception outcomes with linguistic reasoning structures, paving the way toward structural process alignment.
Based on vision tasks, some methods further extend RL to temporal reasoning through audio–visual joint modeling. Video-R1 (Feng et al., 2025) introduces a reward mechanism based on temporal consistency, aiming to enhance the model’s ability to understand the sequential structure of video frames. It compares the predicted frame order with the ground-truth sequence and constructs a reward signal reflecting the logical coherence of temporal reasoning, thereby guiding the model to learn causal inference over time. Although the reward remains output-based and lacks fine-grained feedback over inter-frame relational modeling, the method demonstrates an initial extension of reinforcement learning into multimodal interaction involving vision, audio, and language.
As RL is extended to spatial reasoning, the complexity of reward design increases substantially. In such tasks, models must not only align visual and linguistic modalities semantically, but also reason over action sequences, spatial layouts, and physical constraints within three-dimensional environments. MetaSpatial (Pan and Liu, 2025) systematically constructs a multi-tier reward architecture for 3D scene generation, which includes format validation (e.g., JSON structure correctness, object-label consistency), physics-based verification (e.g., collision detection, boundary adherence), and perceptual evaluation using GPT-4o, which scores generated scenes for visual aesthetics, spatial coherence, and alignment with user preferences. This rich supervision framework supports dense feedback propagation throughout the generation trajectory and reflects the hierarchical dependencies across syntax, physics, and visual perception. By providing full-path feedback across vision, language, and spatial modalities, MetaSpatial establishes a novel and scalable foundation for deploying MLLMs in real-world, physically grounded environments. GFlowVLM (Kang et al., 2025) takes this process-level modeling further by representing reasoning as a non-Markovian decision sequence and adopting Generative Flow Networks (GFlowNets)—a class of generative models that learn to sample structured outputs (e.g., sequences or graphs) with probabilities proportional to their assigned rewards—to model probability distributions over full reasoning trajectories. Through three core loss functions—Trajectory-Balanced, Subtrajectory-Balanced, and Detailed-Balanced—the model ensures both local decisions and global structural coherence are rewarded appropriately. Rather than optimizing isolated outputs, GFlowVLM learns to sample trajectories in proportion to long-term value, enabling a reinforcement framework with structural reward density, consistent feedback across inference stages, and coordinated modality integration.

### v2

@@ 4. RL-Based MLLM Reasoning
R1-Omni (Zhao et al., 2025b) HumanOmni-0.5B SFT&RL GRPO Emotion T
Table 2. Overview of multimodal reasoning models employing reinforcement learning strategies. Univ. Reason denotes general-purpose reasoning across diverse tasks, Image Quality refers to the image quality understanding task, Emotion represents the emotion recognition task and ST-Perception refers to the spatio-temporal perception tasks. In the MM-CoT column, T indicates support for the text modality, I for the image modality.
Deepseek-R1 (Guo et al., 2025) demonstrates that the reasoning capabilities of LLMs can be activated through pure RL. Current RL-based reasoning methods model the reasoning process as a Markov Decision Process (MDP). The core objective is to enhance the model’s reasoning capability by maximizing the expected reward of the reasoning path $S$. Recently, researchers have begun exploring how to apply the R1 training paradigm to MLLMs to further improve their reasoning performance in complex tasks. In Table 2, we summarize the current research progress of reasoning methods based on RL in the multimodal domain. In Section 4.1, we review recent efforts to adapt the R1 training paradigm from LLMs to MLLMs, highlighting key challenges and corresponding solutions. Section 4.2 categorizes reward mechanisms in multimodal settings into Outcome-supervised and Process-supervised types. Section 4.3 discusses methods that enhance training efficiency and stability through data utilization. Finally, Section 4.4 summarizes prevalent RLHF frameworks that support rapid iteration of RL-based reasoning approaches.
@@ 4.1.1. Standardized R1 Training Paradigm in LLM
The success of Kimi K1.5 (Team et al., 2025a) and DeepSeek R1 (Guo et al., 2025) in enhancing model reasoning capabilities through RL has driven the rise of the R1 training paradigm in the field of model reasoning.
@@
Kimi K1.5 (Team et al., 2025a) and Kimi-VL (Team et al., 2025b) employs a variant of the online policy mirror descent (OPMD) algorithm (Mei et al., 2019; Tomar et al., 2020) to apply RL to MLLMs, thereby enhancing their reasoning capabilities. Additionally, it adopts a length penalty mechanism to prevent the model from overthinking. This represents a successful attempt in the direction of MLLM reasoning based on RL. Meanwhile, DeepSeek R1 (Guo et al., 2025) introduces a key innovation by validating that RL with Verifiable Reward (RLVR) is sufficient to effectively elicit reasoning capabilities in LLMs. The core of RLVR lies in its rule-driven reward mechanism, which optimizes model performance efficiently and reliably. Compared to traditional SFT and process supervision methods, RLVR demonstrates significant advantages when handling verifiable tasks, such as mathematical problem-solving and code generation: it eliminates the need for labor-intensive data generation and annotation while significantly improving model performance with limited training data. Additionally, by avoiding the complex training of reward models, RLVR effectively mitigates the issue of reward hacking.
@@
(16) $R(q,o)=\begin{cases}1,&\text{if }o==\text{ground truth},\\ 0,&\text{otherwise}.\end{cases}$
@@
where $R_{format}$ evaluates whether the output of the policy model $\pi_{\theta}$ adheres to predefined template specifications. In DeepSeek R1, the template may require the output to follow the format ¡think¿¡/think¿¡answer¿¡/answer¿, where ¡think¿¡/think¿ contains the model’s reasoning process, and ¡answer¿¡/answer¿ includes the final result. On the other hand, $R_{acc}$ focuses on assessing whether the output in the ¡answer¿¡/answer¿ matches the ground truth. Through this design, RLVR not only ensures the accuracy of the model’s output but also guides the model to generate structured and interpretable reasoning content, thereby enhancing both performance and explainability. In this survey, we refer to the method of using only the format reward and accuracy reward during the RL phase, as defined in Equation 17, as the R1 training paradigm.
@@
While RL has shown promise in improving the reasoning abilities of LLMs, relying solely on large-scale RL is insufficient for robust reasoning development. A fundamental challenge is that reinforcement learning algorithms often require a reasonably good initialization to function effectively. In reasoning tasks, the combinatorial nature of the action space makes it difficult to define intermediate rewards. Consequently, without a meaningful starting policy, exploration becomes inefficient, leading to excessively sparse rewards. This frequently causes training to collapse into trivial or degenerate behaviors. Cold-start methods address this challenge by providing initial scaffolding, including techniques such as curriculum design, synthetic demonstrations, weak supervision, or structured exploration, to bootstrap the model into a regime where RL can be effectively applied. DeepSeek R1 collected thousands of cold-start data to fine-tune the base model, enhancing the readability of reasoning CoT and the potential for RL. Thus, successful reasoning training typically involves a hybrid approach: cold start strategies to seed useful reasoning behaviors, followed by RL to refine and optimize them over time.
@@ 4.1.2. R1 Training Paradim for MLLM
MedVLM-R1 (Pan et al., 2025) extended the DeepSeek R1 training paradigm to visual question-answering (VQA) tasks in the medical domain, guiding the model to answer medical multiple-choice questions through explicit reasoning paths, thereby improving prediction accuracy and OOD generalization performance. Similarly, VisualThinker-R1-Zero (Zhou et al., 2025a) and MM-Eureka (Meng et al., 2025) successfully replicated DeepSeek R1’s long CoT characteristics, including the ”Aha Moment,” self-reflection mechanisms, and dynamic adjustment of response length, in spatial reasoning and mathematical VQA tasks, respectively. These features demonstrate that the experience of DeepSeek R1 can be effectively transferred to complex multimodal reasoning tasks. OThink-MR1 (Liu et al., 2025g) improved the widely used GRPO algorithm by introducing dynamic weighting for the KL divergence term, balancing exploration in the early training stages with exploitation in later stages. To filter data with appropriate difficulty levels for RL training, ThinkLite-VL (Wang et al., 2025e) proposes a sample selection method based on MCTS. This method quantifies the difficulty of each sample by the number of iterations required for the VLM to solve the problem. Despite these efforts achieving the application of the R1 training paradigm in multimodal tasks, they remain constrained by Equation 17 format rewards and accuracy rewards, somewhat neglecting the unique characteristics of tasks in the multimodal domain.
@@
To improve the stability of the RL training process and the upper limit of the model’s inference performance, some methods have begun constructing multi-modal SFT data to cold-start the base model. Vision-R1 (Huang et al., 2025) first leverages VLM and DeepSeek R1 to construct reasoning cold-start data and performs SFT. Subsequently, it employs Progressive Thinking Suppression Training (PTST), a staged RL strategy that incrementally extends the length of CoT reasoning while mitigating overthinking from the early stages. Due to the relative scarcity of image-text data compared to pure text data, LMM-R1 (Peng et al., 2025) proposes a two-stage training strategy to make efficient use of textual CoT data: the first stage involves RL training on pure text data to enhance MLLM’s foundational text reasoning abilities, while the second stage incorporates image-text data to further improve the model’s generalization capabilities. R1-Onevision (Yang et al., 2025) highlights that long reasoning paths may cause the model to overlook visual information. In the SFT stage, R1-Onevision constructs a reasoning dataset incorporating formalized image descriptions. By standardizing the model’s reasoning paths, it enables the model to actively integrate multimodal information during the reasoning process. In the RL stage, the model’s reasoning capabilities and generalization are further strengthened. Skywork R1V (Yi Peng, 2025) proposes a lightweight R1 paradigm training method to alleviate the high consumption of hardware resources. This approach combines a visual encoder with a reasoning-capable LLM backbone through a lightweight visual projection module. The module is iteratively trained using SFT and the R1 paradigm to enhance the reasoning performance of MLLMs. Unlike cold-starting with SFT, Skywork R1V2 (Wei et al., 2025) employs MPO for model cold-starting to address the issue where SFT undermines the performance of subsequent reinforcement learning or reasoning processes. Additionally, a Selective Sample Buffer (SSB) mechanism is designed to address the ”vanishing advantage” problem in GRPO. To address the diversity of mathematical problems, Reason-RFT (Tan et al., 2025) designed customized reward mechanisms for various problems. Each mechanism is tailored to the unique characteristics of its respective problem category, ensuring precise and fair evaluations.
@@
Beyond image-text interactions, some research has explored applying the R1 training paradigm to tasks involving other modalities. R1-Omni (Zhao et al., 2025b) conducted preliminary explorations in mixed-modality tasks involving video and audio. It first performs SFT cold-start with a small amount of reasoning data, then utilized the R1 training paradigm to explore the application of reasoning models in emotion recognition tasks. Spatial-R1 (Ouyang, 2025) extends the R1 training paradigm to the field of video spatial reasoning. Additionally, R1-AQA (Li et al., 2025b), SARI (Wen et al., 2025) and R1-Zero-VSI (Liao et al., 2025) have conducted a series of reinforcement learning-related explorations in audio understanding and video spatial reasoning tasks, respectively. These efforts provide new insights for the further development of multimodal reasoning.
@@ 4.2.1. Outcome-supervised reward Mechanism
Outcome-supervised reward mechanisms represent an advancement in aligning RL with complex, multimodal tasks by leveraging task-specific and cross-modal feedback. These mechanisms not only enhance model performance through structured and hierarchical rewards but also pave the way for more interpretable and generalizable reasoning across diverse modalities. We summarize the reward functions of different reward strategies in Table 3, Table 4, and Table 5, respectively.
@@
Model Reward
Visual-RFT (Liu et al., 2025f)
Detection Tasks: $R=r_{IoU}+r_{conf}+r_{format}$
Classification Task:$R=r_{acc}+r_{format}$
Seg-Zero (Liu et al., 2025e) $R=r_{BIoU}+r_{BL1}+r_{PL1}+r_{SFormat}+r_{format}$
VLM-R1 (Shen et al., 2025) $R_{acc}^{rec}(q,o)=\mathrm{IoU}(b^{*},f_{rec}(o))$, $s_{ovd}=\min(1,\frac{L_{gt}}{L_{pred}})$ $R_{acc}^{ovd}(q,o)=s_{ovd}\cdot\mathbf{mAP}(\mathbf{b}_{pred},\mathbf{b}_{gt})$
Perception-R1 (Yu et al., 2025a) $r_{\mathrm{answer}}=\frac{1}{N}\sum_{i=1}^{N}\Phi\left(y_{i},z_{\hat{\sigma}(i)}\right)$, $R_{\mathrm{total}}=r_{\mathrm{format}}+r_{\mathrm{answer}}$
R1-SGG (Chen et al., 2025c) $R_{\text{total}}=R_{\text{format}}+R_{\text{node}}+R_{\text{edge}}$, $R_{\text{node}}=\lambda_{1}\cdot\text{Sim}(c_{i},\tilde{c}_{j})+\lambda_{2}\cdot\text{IoU}(b_{i},\tilde{b}_{j})+\lambda_{3}\cdot\exp(-\|b_{i}-\tilde{b}_{j}\|_{1})$ $R_{\text{edge}}=\text{Sim}(v_{i},\tilde{v}_{k})\cdot\text{Sim}(v_{j},\tilde{v}_{l})\cdot\text{Sim}(p_{ij},\tilde{p}_{kl})$
VideoChat-R1 (Li et al., 2025c) $R_{\text{st}}=R_{\text{format}}+R_{\text{IoU}}$, $R_{\text{qa}}=R_{\text{format}}+R_{\text{accuracy}}$ $R_{\text{gqa}}=R_{\text{format}}+R_{\text{IoU}}+R_{\text{accuracy}}$, $R_{\text{cap}}=R_{\text{format}}+R_{\text{recall}}$
TinyLLaVA-Video-R1 (Zhang et al., 2025b) $FR=r_{0}+LR$, $LR=\min\left(1,\frac{\text{Len}}{\text{ML}}\right)\cdot r_{1}$ $AR=\begin{cases}r_{2},&\text{if }o\text{ is correct}\\ 0,&\text{otherwise}\end{cases}$ $R=\begin{cases}AR+FR,&\text{if }FR>0\text{ and }AR=r_{2}\\ -FR,&\text{if }FR>0\text{ and }AR=0\\ -(r_{0}+r_{1}+r_{2}),&\text{if }FR=0\end{cases}$
Table 3. Reward function of RL-Based MLLM reasoning methods for Task-Oriented Reward Strategies.
Given the complexity and diversity of multimodal tasks, recent studies have explored reward designs that leverage the intrinsic properties of each modality to broaden the applicability of RL. Visual-RFT (Liu et al., 2025f) targets fundamental computer vision problems such as image classification, object detection, and image-text grounding, introducing task-specific visual rewards. All of these rewards are based on the final prediction accuracy, providing clear label supervision signals within the visual modality. Seg-Zero (Liu et al., 2025e) extends this concept to the field of visual segmentation. In addition to pixel-level IoU and L1 distance metrics used for evaluating segmentation quality, this method introduces model-generated CoT responses as part of the output, enabling partial supervision of the reasoning trajectory. Similarly, VLM-R1 (Shen et al., 2025) designs reward functions for referring expression comprehension and open-vocabulary object detection. To address the multi-agent attributes in visual perception tasks, Perception-R1 (Yu et al., 2025a) formulates the reward matching problem as a bipartite graph matching task and employs the Hungarian algorithm (Kuhn, 1955) to maximize the overall reward. This ensures that each predicted attribute is accurately matched to its corresponding ground truth, thereby optimizing the reward calculation process. R1-SGG (Chen et al., 2025c) and VideoChat-R1 (Li et al., 2025c) both employ modular reward functions, tailored for structured scene graph generation and spatiotemporal video tasks, respectively. They design composite rewards—covering format, node/edge alignment, IoU, and task-specific correctness—to guide model outputs toward structured, temporally grounded responses. In contrast, TinyLLaVA-Video-R1 (Zhang et al., 2025b) applies a lightweight reward scheme combining answer correctness, format adherence, and reasoning length, enabling small models to exhibit interpretable rationales under limited supervision. In addition to designing rewards for single-modality characteristics, some methods have begun to consider cross-modal interactions, representing a transitional phase in the broader shift from single-modality to multi-modality reward modeling.
@@
Model Reward
Q-Insight (Li et al., 2025d) $r^{(i)}=r^{(i)}_{\text{fmt}}+\mathbf{1}_{\text{scr}}\cdot r^{(i)}_{\text{scr}}+\mathbf{1}_{\text{deg}}\cdot\left(\alpha_{1}\cdot r^{(i)}_{\text{deg}}+\alpha_{2}\cdot r^{(i)}_{\text{lev}}\right)$
VLAA-Thinker (Chen et al., 2025b) $R=r_{acc}+\lambda r_{format}$ $R_{open}=1-\exp(-(S_{\theta}(\hat{y})-S_{\theta}(y))\times\beta)\mathrm{~{}if~{}}f_{\theta}(\hat{y})>f_{\theta}(y)\mathrm{~{}else~{}}0$
Relation-R1 (Li et al., 2025a) $r_{\mathrm{form}}(o_{i})=\begin{cases}1&\mathrm{if}~{}o_{i}\text{ adheres to the format},\\ 0&\mathrm{otherwise}.\end{cases}$ $r_{\mathrm{binary}}{(o)}=\alpha\cdot Recall+(1-\alpha)\cdot mRecall$ $r_{\mathrm{n-ary}}(o)=\beta\cdot Acc_{v}+(1-\beta)\cdot Acc_{n}$
FAST (Xiao et al., 2025) $r_{a}=\mathbb{I}(A_{pred}\equiv A_{gt})=\begin{cases}1.0&\mathrm{if~{}}A_{pred}\equiv A_{gt}\\ 0.0&\mathrm{otherwise}\end{cases}$ $r_{f}=\mathbb{I}_{\text{think-tags}}\cdot\mathbb{I}_{\text{answer-tags}}$ $r_{t}=\begin{cases}1-\frac{L}{L_{avg}}&\inf S_{\mathrm{comp.}}<\theta_{comp.}\mathrm{~{}and~{}}r_{a}=1\\ min(\frac{L}{L_{avg}}-1,1)&\inf\theta_{comp.}\leq S_{\mathrm{comp.}}\mathrm{~{}and~{}}r_{a}=0\\ 0&\mathrm{otherwise}\end{cases}$
Table 4. Reward function of RL-Based MLLM reasoning methods for Cross-Modal Interaction Reward Strategies.
To address the non-linear interactions in multimodal reasoning, recent studies assign modality-specific rewards to promote more active cross-modal interactions. Q-Insight (Li et al., 2025d) employs GRPO to jointly optimize score accuracy and distortion classification, enabling perceptual reasoning with minimal supervision and strong zero-shot generalization. Building on this, VLAA-Thinker (Chen et al., 2025b) introduces a mixed reward module combining rule-based and open-ended signals to guide adaptive, realistic reasoning. Relation-R1 (Li et al., 2025a) targets complex visual relations by designing Binary and N-ary Relation Rewards, emphasizing visual-semantic grounding over linguistic bias. To counter overthinking, Fast (Xiao et al., 2025) introduces a complexity-aware thinking reward and adaptively adjusts the KL coefficient based on visual and semantic difficulty. Some methods extend RL to temporal reasoning via audio–visual modeling. Video-R1 (Feng et al., 2025) introduces a reward based on temporal consistency by comparing predicted and ground-truth frame orders, guiding causal inference across time. While the reward remains output-based and coarse-grained, it marks a step toward RL-driven multimodal temporal understanding.
As reinforcement learning extends to spatial reasoning, reward design must accommodate not only semantic alignment across modalities but also reasoning over actions, layouts, and 3D constraints. MetaSpatial (Pan and Liu, 2025) introduces a multi-tier reward system for 3D scene generation, combining format validation, physics-based verification, and GPT-4o-driven perceptual scoring. This hierarchical structure enables dense, cross-modal feedback and supports grounded scene generation. GFlowVLM (Kang et al., 2025) further advances process-level modeling using GFlowNets to sample reasoning trajectories with probabilities aligned to structured rewards. It employs Trajectory-, Subtrajectory-, and Detailed-Balanced losses to optimize coherence at both local and global levels, enabling fine-grained reward propagation throughout multimodal inference.